OpenStudy (anonymous):

how do i find the derivative of: Y=x^3 tan^2 x

myininaya (myininaya):

you have a function times a function so you have to use the product rule

myininaya (myininaya):

so we have y'=[x^3]'(tanx)^2+x^3[(tanx)^2]'

myininaya (myininaya):

(x^3)'=3x^2 using power rule

myininaya (myininaya):

we have [(tanx)^2]'=2(tanx)*[secx]^2 using chain rule

OpenStudy (anonymous):

thank you so much

myininaya (myininaya):

np

OpenStudy (anonymous):

\[3x^2(\tan^2x)+x^3(1/\cos^2x)\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime :)

myininaya (myininaya):

That's not right mmonish91

OpenStudy (anonymous):

we know the equation is uv' + u'v, so take the following: - u = x^3 - u' = 3x^2 - v = tan^2x - v' = 2 tanx sec^2x so: y' = uv' + u'v = (x^3)(2tanxsec^2x) + (3x^2)(tan^2x)

OpenStudy (anonymous):

^_^ simply as that.

OpenStudy (anonymous):

simple*

OpenStudy (anonymous):

@myinninaya please do explain

OpenStudy (anonymous):

mmonish dear, you started correctly, but didn't get d/dx tan^2x right. the derivative of tanx = sec^2x , but in this case the derivative of: (tanx)^ 2 = 2.(tanx)(sec^2x) ; you multiply the power first then subtract it, then you derive tanx. (power rule) ^_^ a^ u = u(a)(u') <-- rule :)

OpenStudy (anonymous):

i see...i left out the 2 for tan^2 mah bad

OpenStudy (anonymous):

a^ u = u(a)^u-1 (u') , sorry lol

OpenStudy (anonymous):

no problem, we all make mistakes ^_^

OpenStudy (anonymous):

i don't really work with sec so my professor said we could leave it in sin or cos format

OpenStudy (anonymous):

but you must know the derivative of tanx = sec^2x while the derivative of secx = secxtanx :)

OpenStudy (anonymous):

which is same as 1/cos^2x

OpenStudy (anonymous):

but I see what you mean :), u left it in cos, it's alright

OpenStudy (anonymous):

yes, true :)

OpenStudy (anonymous):

thank you guys!! ^.^

OpenStudy (anonymous):

you are most welcome sindy ^_^ glad you understood it

OpenStudy (anonymous):

yes i did understand :D ..